If 62.5 moles of an ideal gas is at 6.39 atm at 16.00 °C, what is the volume of the gas?
Answer #1

According to Ideal gas law,

PV = nRT

V = nRT / P

Given

Pressure P = 6.39 atm

No. of moles of gas n = 62.5 moles

Temperature T = 16.00 o C = 16.00 o C + 273 K = 289 K

Universal gas constant R =

Volume V = ?

V = nRT / P

= (62.5 mol * 0.08205 L atm / K mol * 289 K ) / 6.39 atm

= 231.929285 L = 232 L

Volume V = 231.929285 L = 232 L